OpenStudy (anonymous):

In a manufacturing process that laminates several ceramic layers, 1% of the assembles are defective. Assume that the assemblies are independent. a) What is the mean number of assemblies that need to be checked to obtain 5 defective assemblies? b) What is the standard deviation of the number of assemblies that need to be checked to obtain 5 defective assemblies? X = number of assemblies needed to obtain 5 defectives.

OpenStudy (anonymous):

@ybarrap

OpenStudy (ybarrap):

a) If 1% are defective and you want to know how many to check to get 5 defective ones, then 1% of X = 5 or $$ .01X=5\\ \implies X=\cfrac{5}{.01} $$ Round up to closest integer. b) In a discrete uniform distribution, the variance \(\sigma\) is $$ \sigma=\sqrt{\cfrac{n^2-1}{12}} $$ Where n=X found above (must be integer). http://en.wikipedia.org/wiki/Uniform_distribution_%28discrete%29

OpenStudy (ybarrap):

Correction: *In b), the formula I have there is the standard deviation \(\sigma\), not the variance. Standard deviation is the square root of the variance.

OpenStudy (anonymous):

what type of distribution is this : I mean binomial, geometrical Poisson ...

OpenStudy (anonymous):

i guess it should be negative binomial distribution @ybarrap

OpenStudy (ybarrap):

We assume that errors are uniformly distributed. There is no justification for using another type: http://en.wikipedia.org/wiki/Maximum_entropy_probability_distribution#Uniform_and_piecewise_uniform_distributions

OpenStudy (ybarrap):

We assume the defects, not the errors, are uniformly distributed between 1 and X.

OpenStudy (anonymous):

\[σ^{2}=\frac{ k(1-p) }{ p ^{2} }\]

OpenStudy (anonymous):

k=5 p=0.01

OpenStudy (ybarrap):

So you are assuming a geometric distribution. That makes sense.

OpenStudy (anonymous):

yas correct

OpenStudy (ybarrap):

So then the average number of trials for 1st defect = 1/.01 then for 5 defects, it would be X=5*(1/.01)=500.

OpenStudy (ybarrap):

And your variance above would be correct. Just need to take the square root for standard deviation.

OpenStudy (anonymous):

yes , o got it and the standard deviation is the square root of variance

OpenStudy (anonymous):

yes , i did ,,bye the way tnx a lot

OpenStudy (ybarrap):

good question. you're welcome